OpenStudy (anonymous):

giving medal and becoming a fan if somebody can help me with this one question

OpenStudy (anonymous):

Simplify the expression The root of negative nine over the quantity of three minus two i plus the quantity of one plus five i.

OpenStudy (anonymous):

\[\frac{ \sqrt{-9} }{ (3-2i)+(1+5i)}\]

OpenStudy (anonymous):

@whpalmer4 @mathstudent55 @Preetha

OpenStudy (anonymous):

thats not an answer choice

OpenStudy (anonymous):

\[\frac{ 12-9i }{ 25}\]

OpenStudy (anonymous):

\[\frac{ 12-9i }{ 7 }\]

OpenStudy (anonymous):

\[\frac{ 9+12i }{ 25 } \frac{ 9+12i }{ 7 }\]

OpenStudy (whpalmer4):

FIrst thing to do is to simplify the denominator. \[(3−2i)+(1+5i)=\]

OpenStudy (whpalmer4):

\[(3−2i)+(1+5i) = 3-2i+1+5i = 3+1-2i+5i = 4+3i\]Agreed?

OpenStudy (anonymous):

yes i agree

OpenStudy (whpalmer4):

That gives us \[\frac{\sqrt{-9}}{4+3i}\]Now we need to rid the denominator of the complex number. We do that by multiplying both numerator and denominator by the conjugate of the denominator, which is the same number, except we change the sign on the imaginary part. \[\frac{\sqrt{-9}}{(4+3i)}*\frac{(4-3i)}{(4-3i)} = \]Can you do that for me, please?

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

for the denominators do you use the distribute property?

OpenStudy (whpalmer4):

Just multiply \[(4+3i)(4-3i)\]you can use whatever method you use for multiplying polynomials.

OpenStudy (whpalmer4):

or you can use the fact that they are the factors of a difference of squares.

OpenStudy (whpalmer4):

\[a^2-b^2 = (a+b)(a-b)\]

OpenStudy (anonymous):

oh ok

OpenStudy (whpalmer4):

Here it is a number of different ways: FOIL: \[(4+3i)(4-3i) = 4*4 -4*3i + 4*3i -3i*3i = 16 - 12i + 12i -9i^2\]\[\qquad = 16-9i^2\] Distributive property: \[(4+3i)(4-3i) = 4(4-3i) + 3i(4-3i) = 16-12i + 12i - 9i^2\]\[\qquad=16-9i^2\] Difference of squares: \[(4+3i)(4-3i) = (a+b)(a-b) = a^2-b^2 = 4^2-(3i)^2\]\[\qquad=16-9i^2\] In all of those cases, we replace \(i^2 = -1\) to get \[(4+3i)(4-3i) = 16-9i^2 = 16-9(-1) = 16+9\]\[\qquad = 25\] That means our fraction is now \[\frac{\sqrt{-9}(4-3i)}{25}\]

OpenStudy (whpalmer4):

We're getting close to the end :-) \[\sqrt{-9} = \sqrt{-1*9} = \sqrt{-1}*\sqrt{9} = i*3 = 3i\]Right?

OpenStudy (anonymous):

right

OpenStudy (whpalmer4):

Okay, so now we have \[\frac{3i(4-3i)}{25} = \]

OpenStudy (whpalmer4):

You should be able to finish that...

OpenStudy (anonymous):

yes thank you so much:))

OpenStudy (whpalmer4):

you're welcome.